# Exponential Dichotomies in Higher-Dimensional Spatial Dynamics for Elliptic Partial Differential Equations

### v1

@@ 1 Introduction
By expanding the $H^{1}$ norm of $u$, adding and subtracting $r^{2}V_{\infty}$, and applying triangle inequality we obtain
$\frac{1}{r^{2}}\left\|u\right\|_{H^{1}}^{2}=\sum_{k,j}\bigg\|\bigg(\frac{1+\mu_{k}}{r^{2}}\bigg)u_{kj}\bigg\|^{2}_{\mathbb{C}^{d}}$ $\leq\sum_{k,j}\left\|\left(\frac{1-r^{2}V_{\infty}}{r^{2}}+\frac{r^{2}V_{\infty}+\mu_{k}}{r^{2}}\right)u_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}\left(1+\|V_{\infty}\|\right)\left\|u_{kj}\right\|^{2}_{\mathbb{C}^{d}}+\sum_{k,j}\left\|A_{k}(r)^{1/2}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}$
$=C\left\|u\right\|_{L^{2}}^{2}+\left\|A(r)^{1/2}u\right\|_{L^{2}}^{2}.$
If (H2) holds, then we have an $r$-independent lower bound on $|\gamma_{k\ell}(r)|^{2}$ given by two cases:
$|\gamma_{k\ell}(r)|^{2}=|\gamma_{k\ell}(r)^{2}|=\sqrt{({\lambda_{\ell}}^{\mathrm{re}}+\mu_{k}/r^{2})^{2}+\big({\lambda_{\ell}}^{\mathrm{im}}\big)^{2}}\geq\begin{cases}\left|{\lambda_{\ell}}^{\mathrm{im}}\right|&\text{ if }{\lambda_{\ell}}^{\mathrm{im}}\neq 0\\ \,\left|{\lambda_{\ell}}^{\mathrm{re}}\right|&\text{ if }{\lambda_{\ell}}^{\mathrm{im}}=0\end{cases}$ (29)
where we note that the inequality $|\gamma_{k\ell}(r)|^{2}\geq|\lambda_{\ell}^{\mathrm{re}}|$ holds by (H2) only in the case where $\lambda_{\ell}^{\mathrm{im}}=0$.
@@
If $d>1$, then (29) means that $|\gamma_{k\ell}(r)|$ is bounded away from zero uniformly in $k$, $\ell$, and $r$. Let $\Gamma>0$ be this uniform lower bound for all $|\gamma_{k\ell}(r)|$. We can now use a similar argument as that used in Equation (20) to bound $\|u\|_{L^{2}}$ in terms of these scalar coefficients. To do this, recall that $R$ is the matrix whose columns are the eigenvector basis $\{\rho_{1},\ldots,\rho_{d}\}$ for $\mathbb{C}^{d}$ associated with $V_{\infty}$. We derive the desired bound by rewriting $u$ in terms of the eigenvector basis as follows:
$\big\|u\big\|_{L^{2}}^{2}$ $=\sum_{k,j}\|RR^{-1}u_{kj}\|_{\mathbb{C}^{d}}^{2}\leq\sum_{k,j}\|R\|^{2}\big\|R^{-1}u_{kj}\big\|_{\mathbb{C}^{d}}^{2}$
$=\sum_{k,j}\|R\|^{2}\sum_{\ell=1}^{d}|u_{kj\ell}|^{2}\leq\sum_{k,j}\|R\|^{2}\sum_{\ell=1}^{d}\frac{|\gamma_{k\ell}(r)|^{2}}{\Gamma^{2}}|u_{kj\ell}|^{2}$
@@
$=\sum_{k,j}\frac{\|R\|^{2}}{\Gamma^{2}}\left\|R^{-1}A_{k}(r)^{1/2}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}\leq\frac{\|R\|^{2}\|R^{-1}\|^{2}}{\Gamma^{2}}\left\|A(r)^{1/2}u\right\|_{\mathbb{C}^{d}}^{2}.$
@@
$\mathrm{Re}\!\left[\gamma_{k\ell}(r)\right]=\mathrm{Re}\!\left[\left(\lambda_{\ell}+\mu_{k}/r^{2}\right)^{1/2}\right]\geq\mathrm{Re}\!\left[\lambda_{\ell}^{1/2}\right]$
@@
If $\lambda_{\ell}^{\mathrm{im}}\neq 0$, denote the argument of $\gamma_{k\ell}(r)^{2}$ by $\theta_{k\ell}(r)\in[-\pi,\pi)$ so that $\gamma_{k\ell}(r)^{2}=|\gamma_{k\ell}(r)^{2}|e^{i\theta_{k\ell}(r)}$. Similarly, let $\theta_{\ell}$ be the argument of $\lambda_{\ell}$. For the remainder of this proof, assume without loss of generality that $\lambda_{\ell}^{\mathrm{im}}>0$ so that $\theta_{\ell}\in(0,\pi)$. Since $\mu_{k}\geq 0$ and $\gamma_{k\ell}(r)^{2}=\lambda_{\ell}+\mu_{k}/r^{2}$, we have that $\theta_{k\ell}(r)\in(0,\theta_{\ell}]$. Furthermore, since $\mathrm{Im}\left[\gamma_{k\ell}(r)^{2}\right]=\lambda_{\ell}^{\mathrm{im}}$ for all $k$ and $r$, we have that $\left|\gamma_{k\ell}(r)^{2}\right|=\frac{\lambda_{\ell}^{\mathrm{im}}}{\sin(\theta_{k\ell}(r))}$ and we can write the real part of $\gamma_{k\ell}(r)$ directly as a function $A$ of $\theta_{k\ell}(r)$:
$\mathrm{Re}\left[\gamma_{k\ell}(r)\right]=\sqrt{\lambda_{\ell}^{\mathrm{im}}}A(\theta_{k\ell}(r)),\qquad A(\theta):=\frac{\cos(\theta/2)}{\sqrt{\sin(\theta)}},\,\,\theta\in(0,\pi)$
Notice that $\mathrm{Re}[\lambda_{\ell}^{1/2}]=\sqrt{\lambda_{\ell}^{\mathrm{im}}}A(\theta_{\ell})$, and recall that $\theta_{k\ell}(r)\in(0,\theta_{\ell})$. To show that $\mathrm{Re}\left[\gamma_{k\ell}(r)\right]\geq\mathrm{Re}[\lambda_{\ell}^{1/2}]$, it is therefore sufficient to show that the function $A(\theta)$ is decreasing on $(0,\theta_{\ell}]$.
Taking a derivative, applying product rule and simplifying with double-angle trigonometric identities,
$A^{\prime}(\theta)$ $=\frac{-\cos(\theta/2)\cos(\theta)}{2\sin^{3/2}(\theta)}+\frac{-\sin(\theta/2)}{2\sin^{1/2}(\theta)}=\frac{-\cos(\theta/2)\left[\cos(\theta)+2\sin^{2}(\theta/2)\right]}{2\sin^{3/2}(\theta)}=\frac{-\cos(\theta/2)}{2\sin^{3/2}(\theta)}$
Since both $\cos(\theta/2)$ and $\sin^{3/2}(\theta)$ are strictly positive for $\theta\in(0,\pi)$, $A^{\prime}(\theta)$ is strictly negative and $A$ is a decreasing function on $\theta\in(0,\pi)$. It follows that
$\mathrm{Re}\left[\gamma_{k\ell}(r)\right]=\sqrt{\lambda_{\ell}^{\mathrm{im}}}A(\theta_{k\ell}(r))\geq\sqrt{\lambda_{\ell}^{\mathrm{im}}}A(\theta_{\ell})=\mathrm{Re}\left[\lambda_{\ell}^{1/2}\right]\quad\text{ for all }k>0\text{ and }r>1.$

### v2

@@ 1 Introduction
By expanding the $H^{1}$ norm of $u$, adding and subtracting $r^{2}V_{\infty}$, and applying both triangle and Young’s inequalities we obtain
$\frac{1}{r^{2}}\left\|u\right\|_{H^{1}}^{2}$ $=\sum_{k,j}\bigg\|\bigg(\frac{1+\mu_{k}}{r^{2}}\bigg)^{1/2}u_{kj}\bigg\|^{2}_{\mathbb{C}^{d}}\leq\sum_{k,j}\left\|\left(\left(1+\mu_{k}/r^{2}\right)^{1/2}\pm A_{k}(r)^{1/2}\right)u_{kj}\right\|_{\mathbb{C}^{d}}^{2}$
$\leq\sum_{k,j}\Bigg[\left\|\left(\left(1+\mu_{k}/r^{2}\right)^{1/2}-A_{k}(r)^{1/2}\right)u_{kj}\right\|_{\mathbb{C}^{d}}+\left\|A_{k}(r)^{1/2}u_{kj}\right\|_{\mathbb{C}^{d}}\Bigg]^{2}$
$\leq 2\bigg[\sum_{k,j}\left\|(1+\mu_{k}/r^{2})^{1/2}-A_{k}(r)^{1/2}\right\|^{2}_{L(\mathbb{C}^{d})}\left\|u_{kj}\right\|_{\mathbb{C}^{d}}^{2}\bigg]+2\left\|A(r)^{1/2}u\right\|_{L^{2}}^{2}$
We proceed by showing that the matrix norm of $(1+\mu_{k}/r^{2})^{1/2}-A_{k}(r)^{1/2}$ is bounded above by a finite constant $C\geq 0$. Rewriting the matrix norm, we get
$\left\|(1+\mu_{k}/r^{2})^{1/2}-A_{k}(r)^{1/2}\right\|$ $=\left\|\frac{1+\mu_{k}/r^{2}-V_{\infty}-\mu_{k}/r^{2}}{(1+\mu_{k}/r^{2})^{1/2}+A_{k}(r)^{1/2}}\right\|=\frac{\left\|1-V_{\infty}\right\|}{\left\|(1+\mu_{k}/r^{2})^{1/2}+A_{k}(r)^{1/2}\right\|}.$
Finally, since $\mu_{k}/r^{2}\geq 0$ the denominator has a norm of at least $1$, and we have
$\left\|(1+\mu_{k}/r^{2})^{1/2}-A_{k}(r)^{1/2}\right\|$ $\leq\|1-V_{\infty}\|.$
Setting $C=\|1-V_{\infty}\|\geq 0$, the first statement follows.
To prove the second statement, notice that if (H2) holds, then we have an $r$-independent lower bound on $|\gamma_{k\ell}(r)|^{2}$ given by two cases:
$|\gamma_{k\ell}(r)|^{2}=|\gamma_{k\ell}(r)^{2}|=\sqrt{\left(\mathrm{Re}\left[{\lambda_{\ell}}\right]+\mu_{k}/r^{2}\right)^{2}+\mathrm{Im}\left[{\lambda_{\ell}}\right]^{2}}\geq\begin{cases}\big|\mathrm{Im}\left[{\lambda_{\ell}}\right]\big|&\text{ if }\mathrm{Im}\left[{\lambda_{\ell}}\right]\neq 0\\ \,\big|\mathrm{Re}\left[{\lambda_{\ell}}\right]\big|&\text{ if }\mathrm{Im}\left[{\lambda_{\ell}}\right]=0\end{cases}$ (24)
where we note that the inequality $\big|\gamma_{k\ell}(r)\big|^{2}\geq\big|\mathrm{Re}\left[\lambda_{\ell}\right]\big|$ holds by (H2) only in the case where $\mathrm{Im}\left[\lambda_{\ell}\right]=0$.
@@
If $d>1$, then (24) means that $|\gamma_{k\ell}(r)|$ is bounded away from zero uniformly in $k$, $\ell$, and $r$. Let $\Gamma>0$ be this uniform lower bound for all $|\gamma_{k\ell}(r)|$, and note that $1\leq\frac{|\gamma_{k\ell}(r)|}{\Gamma}$. Using this bound, and the decomposition of the $L^{2}$ norm given in (3.1), we have
$\big\|u\big\|_{L^{2}}^{2}$ $=\sum_{k,j}\|u_{kj}\|_{\mathbb{C}^{d}}^{2}=\sum_{k,j}\|RR^{-1}u_{kj}\|_{\mathbb{C}^{d}}^{2}\leq\sum_{k,j}\|R\|^{2}\big\|R^{-1}u_{kj}\big\|_{\mathbb{C}^{d}}^{2}$
where $R$ is the matrix whose columns are the eigenvectors $\{\rho_{1},\ldots,\rho_{d}\}$ of $V_{\infty}$, guaranteed by (H3). Let $u_{kj}=\sum_{\ell=1}^{d}u_{kj}^{(\ell)}\rho_{\ell}$. Then $R^{-1}u_{kj}=(u_{kj}^{(1)},\ldots,u_{kj}^{(d)})\in\mathbb{C}^{d}$, and we have
$\|R^{-1}u_{kj}\|^{2}_{\mathbb{C}^{d}}=\sum_{\ell=1}^{d}\left|u_{kj}^{(\ell)}\right|^{2}.$
We proceed by bounding $\|u\|_{L^{2}}^{2}$ as follows:
$\big\|u\big\|_{L^{2}}^{2}$ $=\sum_{k,j}\|R\|^{2}\sum_{\ell=1}^{d}\left|u_{kj}^{(\ell)}\right|^{2}$
$\leq\sum_{k,j}\|R\|^{2}\sum_{\ell=1}^{d}\frac{|\gamma_{k\ell}(r)|^{2}}{\Gamma^{2}}\left|u_{kj}^{(\ell)}\right|^{2}$
@@
Finally, recall that by the diagonalization of $V_{\infty}$, (H3), we have that $R^{-1}A_{k}(r)^{1/2}=\,\,\mathrm{diag}(\gamma_{k\ell}(r))R^{-1}$. Therefore,
$\big\|u\big\|_{L^{2}}^{2}$ $\leq\sum_{k,j}\frac{\|R\|^{2}}{\Gamma^{2}}\left\|R^{-1}A_{k}(r)^{1/2}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}\leq\frac{\|R\|^{2}\|R^{-1}\|^{2}}{\Gamma^{2}}\left\|A(r)^{1/2}u\right\|_{\mathbb{C}^{d}}^{2}.$
@@
$\mathrm{Re}\Big[\gamma_{k\ell}(r)\Big]=\mathrm{Re}\!\left[\left(\lambda_{\ell}+\mu_{k}/r^{2}\right)^{1/2}\right]\geq\mathrm{Re}\!\left[\lambda_{\ell}^{1/2}\right]$
@@
If $\mathrm{Im}\left[\lambda_{\ell}\right]\neq 0$, then standard calculation of a complex principal square root gives
$\mathrm{Re}\!\left[\left(\lambda_{\ell}+\mu_{k}/r^{2}\right)^{1/2}\right]^{2}=\frac{1}{2}\Big(\left|\lambda_{\ell}+\mu_{k}/r^{2}\right|+\mathrm{Re}\!\left[\lambda_{\ell}+\mu_{k}/r^{2}\right]\Big)=\frac{g_{\ell}(\mu_{k}/r^{2})}{2}$
where we define
$g_{\ell}(\xi):=\left[\left(\mathrm{Re}\left[\lambda_{\ell}\right]+\xi\right)^{2}+\mathrm{Im}\left[\lambda_{\ell}\right]^{2}\right]^{1/2}+(\mathrm{Re}\left[\lambda_{\ell}\right]+\xi).$
A straightforward calculation shows that $g_{\ell}(\xi)$ is strictly increasing for all $\xi\geq 0$ with $\mathrm{Im}\left[\lambda_{\ell}\right]\neq 0$. Therefore we have
$\mathrm{Re}\!\left[\left(\lambda_{\ell}+\mu_{k}/r^{2}\right)^{1/2}\right]^{2}=\frac{g_{\ell}(\mu_{k}/r^{2})}{2}\geq\frac{g_{\ell}(0)}{2}=\frac{1}{2}\Big(\left|\lambda_{\ell}\right|+\mathrm{Re}\left[\lambda_{\ell}\right]\Big)=\mathrm{Re}\!\left[\lambda_{\ell}^{1/2}\right]^{2}.$
The result follows since we are taking the principal square root. ∎